OpenStudy (anonymous):

how to solve 4x - (x + 1) = 2x + 4

OpenStudy (rushwr):

\[4x-x-1 =2x +4\] \[3x -1 = 2x+4\] \[3x -2x = 5\] \[x = 5\]

OpenStudy (anonymous):

can you explain it plz

OpenStudy (rushwr):

get rid of the brackets first for that we have multiply the stuff inside the brackets by the negative mark